OpenStudy (anonymous):

cathrynmolle14 Medals 0 I have 2 questions I need help with: 1) what is the slope of a line that passes through the point (-1,1) and is parallel to a like that passes through (4,6) and (-1,-4) 2) Blake has a total of 64 DVD's and CD's. If the number of the CD's is 8 more times the number of DVDs how many DVD's does he have? PLEASE I NEED HELP!

OpenStudy (jdoe0001):

parallel lines have the same slope, if the line at point "(-1,1) and is parallel to a line that passes through (4,6) and (-1,-4)" then the slope of the 2nd line at (4,6) and (-1,-4)" is the same as the line at (-1, 1), that is \(\bf \begin{array}{lllll} &x_1&y_1&x_2&y_2\\ &(4\quad ,&6)\quad &(-1\quad ,&-4) \end{array} \\\quad \\ slope = m= \cfrac{rise}{run} \implies \cfrac{y_2-y_1}{x_2-x_1}\) once you find the slope \(\bf \begin{array}{lllll} &x_1&y_1\\ &(-1\quad ,&1) \end{array} \\\quad \\ slope = m= \square ? \\ \quad \\ y-y_1=m(x-x_1)\qquad \textit{plug in your values and solve for "y"}\) ------------------------------------------------- there are 64 Dvds and Cds, so say Dvd = d Cd = c \(\bf d + c = 64\) "CD's is 8 more times the number of DVDs " so whatever Dvds amount is, "d", then Cds are 8 times that much, that is, "8d" \(\bf d +c = 64\\ c = 8d\quad thus\\ d + 8d = 64\) solve for "d" to get the Dvds amount

OpenStudy (anonymous):

for problem 2i does D=8?

OpenStudy (anonymous):

thank you for your help!